A large grain silo consists of a cylinder measuring 30 feet tall, a diameter at the base of 18 feet, and a hemisphere (half-dome) on top. How many gallons of paint will be required to paint this silo if a gallon of the paint to be used covers 500 square feet?
Answer by lwsshak3(11628)

A large grain silo consists of a cylinder measuring 30 feet tall, a diameter at the base of 18 feet, and a hemisphere (half-dome) on top. How many gallons of paint will be required to paint this silo if a gallon of the paint to be used covers 500 square feet.
...
formulas used:
Surface area of sphere=4πr^2
Circumference = 2πr
r=diameter/2=18/2=9 ft
height, h=30 ft
..
Surface area of half-dome=1/2*4*π*r^2=2π*81=162π
Surface area of bottom=π*r^2=81π
Surface area of cylinder=circumference*height=2πr*30=540π
Total surface area of silo=(162+81+540)π=783π sq ft
ans:
Paint required=783π/500=4.92gal